# Charge carrier relaxation dynamics in the one-dimensional Kondo lattice model

### v1

@@ I Introduction
Theoretical research into the nonequilibrium dynamics of strongly-correlated quantum many-body systems continues to be fueled from several directions. First, the experimental advances with ultrafast dynamics provide access to ever-more time windows and observables, unveiling physics that has been out of reach so far [1, 2, 3, 4, 5, 6, 7, 8, 9]. This includes the short-time dynamics in the electronic relaxation dynamics [9, 5], stabilization of metastable states [2, 8], and controlled driving of materials into states with a desired type of order [4, 7]. Second, modeling the relevant materials often requires accounting for several degrees of freedom, such as itinerant electrons, orbital degrees of freedom, phonons, or magnetic excitations, giving rise to a high degree of complexity [10, 11]. Third, theoretical descriptions face formidable challenges, since several decades in time need to be covered with different physics dominating different time windows [12, 13, 14]. This is usually not possible without approximations, and therefore, simple but generic model Hamiltonians such as the Hubbard model [15] or Kondo-lattice models [16, 17, 18] continue to play an important role.
@@
In this work, we will take the route of working with a model system, and we are specifically interested in the relaxation dynamics of electrons coupled to a magnetic subsystem. This falls into a broader class of problems, where one seeks to unravel the electronic time evolution after an optical excitation due to a coupling to bosonic degrees of freedom [19, 20, 21, 22]. The latter are usually either phonons or magnons. Several generic initial conditions have been studied: (i) the relaxation of a single charge carrier with a well-defined quasi-momentum in an empty band [22, 21, 23], (ii) the dynamics of a charge carrier in real-space [21, 22, 24], (iii) the relaxation of one or many electrons after an optical excitation modeled as a pulse [8, 20, 15], (iv) the specific analysis of the time evolution of an order parameter $\mathcal{O}(t)$ after such initial conditions, either concerning their decay as a function of time [25, 26, 19, 27, 28] or the emergence of a non-zero $\mathcal{O}(t)$ due to the optical excitation [29, 30, 31].
@@
In our work, we start from the first setup, which we will modify as outlined below. We pursue the strategy of obtaining numerically exact results, and therefore, we restrict the discussion to the one-dimensional case. As a model Hamiltonian, we consider a generalized Kondo-lattice Hamiltonian [32, 33, 34, 35, 36, 37, 38, 39] as sketched in Fig. 1(a), incorporating the Kondo coupling between localized spins and the itinerant electrons, a direct coupling between the spin degrees of freedom, and disorder. On the one hand, we aim at establishing a comprehensive understanding of the relaxation dynamics depending on the model parameters and for idealized initial conditions in the spin sector. On the other hand, we want to address the issue of thermalization.
@@
Both the cases of a coupling of an optically excited electron to phonons and magnons have been investigated. For the former, we mention [22, 40, 41, 42], while for the latter, Refs. [24, 43, 44] are examples. As an ultimate goal beyond the scope of the present study, we wish to understand the competition between several bosonic relaxation channels, e.g., in the presence of phonons and magnons (see [45, 46] for work in that direction).
@@
To set the stage, we paraphrase results from Ref. [21], where a charge carrier was initialized in a quasi-momentum eigenstate in an empty band and then coupled to optical phonons via a Holstein coupling. In the adiabatic regime where phonon frequencies $\omega_{0}$ are smaller than the electronic bandwidth, Ref. [21] reports a fast relaxation to a quasi-stationary state. That quasi-stationary state is characterized by a complete energy transfer from the electron to the optical phonons accompanied by a relaxation of the average electronic momentum to zero. Interestingly, the final momentum distribution appears to be thermal already in a simple model such as the Holstein chain studied in [21] (substantiated by follow-up studies [47, 48]). Moreover, the typical time scale for a complete momentum and energy transfer was extracted in dependence of relevant parameters such as the electron-phonon coupling and the oscillator frequency. In the present study, we aim at establishing similarities and differences between the relaxation dynamics in the Holstein model and the Kondo-lattice chain.
@@
As a starting point, let us consider the well-studied case (see, e.g., [49, 50, 51, 52, 53, 54]) of a single charge carrier moving in the ferromagnetic background (short FM) of the localized spins [see Fig. 1(a)], thus an initial state of the form:
@@
In this exactly solvable case [50, 51, 52] of effectively two particles – electron and one emitted magnon – the system relaxes, yet evidently not to a thermal state, as evidenced by the persistence of finite-momentum peaks in the quasi-momentum distribution. Likewise, for the case of two charge carriers in a ferromagnetic background, no thermalization is observed as these are still few-body problems in a subspace with subexponential scaling with system size [55, 56, 57]. We will start by illustrating these behaviors using numerical simulations to provide a comprehensive basis for our study. We further consider the addition of diagonal disorder seen by the conduction electron. As expected, Anderson localization leads to an additional suppression of thermalization. However, at weak disorder, we still observe a complete magnetization transfer. For large disorder and a large $J$, the dynamics is purely oscillatory, dominated by local dimers.
@@ II.3 Numerical Method
We use the time-dependent Lanczos method [58] to simulate the time evolution of our many-body system and full exact diagonalization to compute thermal expectation values [59, 60]. We use periodic boundary conditions in all simulations.
We exploit symmetries, including total magnetization $S^{z}_{\text{total}}$, conservation of the number of conductions electrons $N_{e}$, and translational invariance [60]. The latter symmetry arises from the invariance of the Hamiltonian under lattice translations, i.e., eigenstates of $\hat{H}$ can also be chosen as eigenstates of the lattice translation operator $\hat{T}$, with eigenvalues $e^{iK}$, where the crystal momentum $K=2\pi n/L$ ($n=0,1,\ldots,L-1$) labels the different momentum sectors.
@@
For the propagation by one time step, we start by finding the Krylov subspace by iteratively applying the Hamiltonian to the current state $\left|\Psi(t)\right$> $D_{K}-1$ times, and construct the operator $T_{D}$ in the resulting Krylov basis. We find that $D_{K}=20$ is usually sufficient, as we verified by varying $D_{K}$. To diagonalize the matrix, we use the LAPACK package [61] and a self-written code for the matrix multiplication to get the time-evolved state $\left|\Psi(t+\Delta t)\right>=V^{\dagger}_{D}e^{-i\Delta tT_{D}}V_{D}\left|\Psi(t)\right$>. Next, we use the quantum state $\left|\Psi(t+\Delta t)\right$> to compute expectation values of operators at $t+\Delta t$ time. Additional parameters used in this method are the time step, which has a range between $\Delta t/t_{0}=0.005$ and $\Delta t/t_{0}=0.05$ (depending on the regime and initial conditions), and a maximum system size of $L=226$ for the initial state $|\Psi_{0,F}\rangle$.
@@ II.4 Thermal expectation values
Assuming thermalization [62, 63], the expected temperature is obtained by equating the energy in the initial state to the thermal expectation value:
@@ II.5 Short-time evolution
It is instructive to determine the short-time evolution analytically. To that end, we expand the time-evolution operator $\hat{U}(t)=\mbox{exp}(-i\hat{H}t)$. However, our Hamiltonian is composed of non-commuting terms. Using the Baker-Campbell-Hausdorff (BCH) formula [64], we get the relation
@@ III Relaxation dynamics of a single charge carrier in a ferromagnetic background
The case without disorder has been solved exactly and was investigated from many perspectives in the literature [50, 51, 52]. We provide a discussion of this case to set up the study of the many-body case, putting the focus on a qualitative picture and the behavior of the quasi-momentum distribution function.
In [52] (extending earlier work, see references in [52]) all eigenstates on finite systems were constructed. The following results were obtained for the time-dependence of $S_{c}^{z}$: (i) the short-time dynamics of the spin polarization is quadratic in time (which is generic), (ii) in an intermediate time-window, the decay is exponential, (iii) the long-time behavior deviates from an exponential decay due to the van-Hove singularity at the bottom of the band, and (iv) oscillations occur at large $J/t_{0}$, The oscillations can be interpreted as an indication of spin-polaron formation (see, e.g., [53]).
Several aspects of the relaxation dynamics in the regime of $J\ll t_{0}$ can be understood by considering the relaxation of a two-level system coupled to a generic bosonic bath, for which there is a substantial amount of literature [65, 66, 67, 68]. Essentially, the conduction electron’s polarization is transferred in a single spin-flip process by creating an excitation in the spin bath, thus preventing the electronic kinetic energy from further relaxation. After a short coherent dynamics, there is a Markovian regime with an exponential decay with a time scale $J^{2}/t_{0}$. For $J/t_{0}\lesssim 1.5$, we find an excellent agreement between numerical data and these approximate analytical results.
@@ III.1 Numerical illustration of relaxation dynamics
The initial state with a complete parallel alignment of all localized spins and antiparallel spin orientation of the conduction electron belongs to the subspace $S^{z}_{\text{total}}=\frac{1}{2}(L-1)$. In principle, this coincides with the total spin quantum number of the ground state of the Kondo-lattice model with one itinerant electron. However, in the ground state, the lowest energy is obtained in the momentum sector $K=0$ when the conduction electron has the momentum $k_{0}=0$, implying that our initial state is well-separated from the ground state [50]. For small $J/t_{0}$ (in the range $0.1<J/t_{0}<2.0$), the initial state becomes essentially orthogonal to the ground state, with its full spectral weight residing in the excited states, as illustrated in App. A.
@@ III.1.1 Regime of $J\ll 2t_{0}$
Based on these observations, we conclude that, in the limit of $J\ll 2t_{0}$, the conduction electron undergoes a scattering process with four components: the electron acquires a backward-propagating ($bp$) and a forward-propagating ($fp$) component, and the same for the emitted magnon [69, 70]. After this scattering process, the $z$-component of the electron spin is spin-down and $k_{bp}=-k_{fp}$. Due to momentum conservation, one magnon is excited with weight at the opposite momenta $q^{m}_{bp}=-q^{m}_{fp}$ (up to umklapp scattering).
@@ III.1.2 Two-level picture for $J/t_{0}\ll 1$
For such a situation, standard approximations (see App. B) yield a Markovian behavior and hence an exponential decay is found [71], in agreement with the exact solution [52]. We extract the decay constant from a fit to the numerical data, resulting in these functional dependencies for the expectation values:
@@
In Fig. 2, we compare the computational results (solid lines) with the TLS results (dashed lines), where one can observe a good agreement for all operators up to $J/t_{0}\lesssim 1.5$. Note that as $J/t_{0}$ increases, oscillations appear. These correspond to magnon reabsorptions and reemissions not captured in the Markovian approximation. Further, the very short-time window is not captured either since there, non-Markovian processes are always present (see, e.g., [72, 73]).
@@ III.1.3 Energy transfer and crossover regime $J\sim 2t_{0}$
For small $J/t_{0}$, a complete magnetization transfer takes place on time scales proportional to $4t_{0}/J^{2}$. At the same time, the kinetic energy of the electron cannot be further reduced after this process since the energy transfer is tied to magnetization transfer. In the small $J/t_{0}$ regime, the energy transferred $\Delta E$ to the spin system is proportional to the exchange coupling. However, as shown in Fig. 5, this linear dependence breaks down when $J/t_{0}\gtrsim 1$. In this intermediate coupling regime, the momentum transfer approaches $\Delta k\approx\pi/2$ as $J/t_{0}\approx 2$. Thus, the initial state couples to the region of the maximum density of states in the dispersion. Therefore, as mentioned before, additional effects beyond the simple TLS picture appear, such as repeated absorption and emission of the magnon, the van-Hove singularity at the band bottom becomes increasingly relevant [52], and the overlap of the initial state becomes significant with more than two many-body eigenstates. Consequently, the TLS description breaks down and fails to capture the relaxation dynamics for $J\gtrsim t_{0}$.
@@ III.1.4 Regime of $J\gg 2t_{0}$
with fitting parameters $a=0.75$ and $b=0.72$. The Gaussian decay indicates a non-Markovian dynamics [73, 74], which is represented as a black dashed line in Fig. 6. Such a Gaussian decay is also observed in central spin models, where it arises due to a contribution of many individual local environments of localized spins seeing a fluctuating mean-field due the interaction with the central spin [75, 76, 74, 77].
In the long-time regime ($t\gg 1/t_{0}$), shown in the inset of Fig. 6 using a log-log scale, the envelope follows a straight line (red dashed line), suggesting a power-law decay proportional to $t^{-3/2}$ due to the van-Hove singularity at the band bottom, in agreement with the exact analytical solution [52].

### v2

@@ I Introduction
Theoretical research into the nonequilibrium dynamics of strongly-correlated quantum many-body systems continues to be fueled from several directions. First, the experimental advances with ultrafast dynamics provide access to ever-more time windows and observables, unveiling physics that has been out of reach so far [boschini2024, zhang2016, Giannetti2016, mitrano2016, dal2015, orenstein2012, fausti2011, yusupov2010, cavalleri2001]. This includes the short-time dynamics in the electronic relaxation dynamics [cavalleri2001, dal2015], stabilization of metastable states [zhang2016, yusupov2010], and controlled driving of materials into states with a desired type of order [mitrano2016, fausti2011]. Second, modeling the relevant materials often requires accounting for several degrees of freedom, such as itinerant electrons, orbital degrees of freedom, phonons, or magnetic excitations, giving rise to a high degree of complexity [basov2017, dagotto2005]. Third, theoretical descriptions face formidable challenges, since several decades in time need to be covered with different physics dominating different time windows [wilner2014, aoki2014, Moshnyaga2025]. This is usually not possible without approximations, and therefore, simple but generic model Hamiltonians such as the Hubbard model [Murakami2025] or Kondo-lattice models [Werner2012, zhu2021, chen2024] continue to play an important role.
@@
In this work, we will take the route of working with a model system, and we are specifically interested in the relaxation dynamics of electrons coupled to a magnetic subsystem. This falls into a broader class of problems, where one seeks to unravel the electronic time evolution after an optical excitation due to a coupling to bosonic degrees of freedom [Stolpp2020, Koehler2018, Dorfner2015, Golez2012]. The latter are usually either phonons or magnons. Several generic initial conditions have been studied: (i) the relaxation of a single charge carrier with a well-defined quasi-momentum in an empty band [Golez2012, Dorfner2015, Lenarcic2014], (ii) the dynamics of a charge carrier in real-space [Dorfner2015, Golez2012, Mierzejewski2011], (iii) the relaxation of one or many electrons after an optical excitation modeled as a pulse [yusupov2010, Koehler2018, Murakami2025], (iv) the specific analysis of the time evolution of an order parameter $\mathcal{O}(t)$ after such initial conditions, either concerning their decay as a function of time [polkovnikov2011, Fauseweh2020, Stolpp2020, osterkorn22, Paprotzki2024] or the emergence of a non-zero $\mathcal{O}(t)$ due to the optical excitation [Shirakawa2020, Paeckel2020, Fauseweh2025].
@@
In our work, we start from the first setup, which we will modify as outlined below. We pursue the strategy of obtaining numerically exact results, and therefore, we restrict the discussion to the one-dimensional case. As a model Hamiltonian, we consider a generalized Kondo-lattice Hamiltonian [ueda1991, tsunetsugu93, Tsunetsugu1997, gulacsi06, trebst2006, basylko08, smerat09, arredondo12] as sketched in Fig. 1(a), incorporating the Kondo coupling between localized spins and the itinerant electrons, a direct coupling between the spin degrees of freedom, and disorder. On the one hand, we aim at establishing a comprehensive understanding of the relaxation dynamics depending on the model parameters and for idealized initial conditions in the spin sector. On the other hand, we want to address the issue of thermalization.
@@
Both the cases of a coupling of an optically excited electron to phonons and magnons have been investigated. For the former, we mention [Golez2012, sentef2016, park2022, caruso2022], while for the latter, Refs. [Mierzejewski2011, Golez2014, mondal2019] are examples. As an ultimate goal beyond the scope of the present study, we wish to understand the competition between several bosonic relaxation channels, e.g., in the presence of phonons and magnons (see [Kogoj2014, Kogoj2016] for work in that direction).
@@
To set the stage, we paraphrase results from Ref. [Dorfner2015], where a charge carrier was initialized in a quasi-momentum eigenstate in an empty band and then coupled to optical phonons via a Holstein coupling. In the adiabatic regime where phonon frequencies $\omega_{0}$ are smaller than the electronic bandwidth, Ref. [Dorfner2015] reports a fast relaxation to a quasi-stationary state. That quasi-stationary state is characterized by a complete energy transfer from the electron to the optical phonons accompanied by a relaxation of the average electronic momentum to zero. Interestingly, the final momentum distribution appears to be thermal already in a simple model such as the Holstein chain studied in [Dorfner2015] (substantiated by follow-up studies [Jansen2019, Schoenle2021]). Moreover, the typical time scale for a complete momentum and energy transfer was extracted in dependence of relevant parameters such as the electron-phonon coupling and the oscillator frequency. In the present study, we aim at establishing similarities and differences between the relaxation dynamics in the Holstein model and the Kondo-lattice chain.
@@
As a starting point, let us consider the well-studied case (see, e.g., [Shastry81, sigrist91, Nakano2012, Henning2012, Moeller2013, Frakulla2024]) of a single charge carrier moving in the ferromagnetic background (short FM) of the localized spins [see Fig. 1(a)], thus an initial state of the form:
@@
In this exactly solvable case [sigrist91, Nakano2012, Henning2012] of effectively two particles – electron and one emitted magnon – the system relaxes, yet evidently not to a thermal state, as evidenced by the persistence of finite-momentum peaks in the quasi-momentum distribution. Likewise, for the case of two charge carriers in a ferromagnetic background, no thermalization is observed as these are still few-body problems in a subspace with subexponential scaling with system size [Moeller2012a, Moeller2012, Rausch2019]. We will start by illustrating these behaviors using numerical simulations to provide a comprehensive basis for our study. We further consider the addition of diagonal disorder seen by the conduction electron. As expected, Anderson localization leads to an additional suppression of thermalization. However, at weak disorder, we still observe a complete magnetization transfer. For large disorder and a large $J$, the dynamics is purely oscillatory, dominated by local dimers.
@@ II.3 Numerical Method
We use the time-dependent Lanczos method [Manmana2005] to simulate the time evolution of our many-body system and full exact diagonalization to compute thermal expectation values [fehske2007, sandvik2010]. We use periodic boundary conditions in all simulations.
We exploit symmetries, including total magnetization $S^{z}_{\text{total}}$, conservation of the number of conductions electrons $N_{e}$, and translational invariance [sandvik2010]. The latter symmetry arises from the invariance of the Hamiltonian under lattice translations, i.e., eigenstates of $\hat{H}$ can also be chosen as eigenstates of the lattice translation operator $\hat{T}$, with eigenvalues $e^{iK}$, where the crystal momentum $K=2\pi n/L$ ($n=0,1,\ldots,L-1$) labels the different momentum sectors.
@@
For the propagation by one time step, we start by finding the Krylov subspace by iteratively applying the Hamiltonian to the current state $\left|\Psi(t)\right$> $D_{K}-1$ times, and construct the operator $T_{D}$ in the resulting Krylov basis. We find that $D_{K}=20$ is usually sufficient, as we verified by varying $D_{K}$. To diagonalize the matrix, we use the LAPACK package [anderson1999] and a self-written code for the matrix multiplication to get the time-evolved state $\left|\Psi(t+\Delta t)\right>=V^{\dagger}_{D}e^{-i\Delta tT_{D}}V_{D}\left|\Psi(t)\right$>. Next, we use the quantum state $\left|\Psi(t+\Delta t)\right$> to compute expectation values of operators at $t+\Delta t$ time. Additional parameters used in this method are the time step, which has a range between $\Delta t/t_{0}=0.005$ and $\Delta t/t_{0}=0.05$ (depending on the regime and initial conditions), and a maximum system size of $L=226$ for the initial state $|\Psi_{0,F}\rangle$.
@@ II.4 Thermal expectation values
Assuming thermalization [rigol08, DAlessio2016], the expected temperature is obtained by equating the energy in the initial state to the thermal expectation value:
@@ II.5 Short-time evolution
It is instructive to determine the short-time evolution analytically. To that end, we expand the time-evolution operator $\hat{U}(t)=\mbox{exp}(-i\hat{H}t)$. However, our Hamiltonian is composed of non-commuting terms. Using the Baker-Campbell-Hausdorff (BCH) formula [varadarajan2013], we get the relation
@@ III Relaxation dynamics of a single charge carrier in a ferromagnetic background
The case without disorder has been solved exactly and was investigated from many perspectives in the literature [sigrist91, Nakano2012, Henning2012]. We provide a discussion of this case to set up the study of the many-body case, putting the focus on a qualitative picture and the behavior of the quasi-momentum distribution function.
In [Henning2012] (extending earlier work, see references in [Henning2012]) all eigenstates on finite systems were constructed. The following results were obtained for the time-dependence of $S_{c}^{z}$: (i) the short-time dynamics of the spin polarization is quadratic in time (which is generic), (ii) in an intermediate time-window, the decay is exponential, (iii) the long-time behavior deviates from an exponential decay due to the van-Hove singularity at the bottom of the band, and (iv) oscillations occur at large $J/t_{0}$, The oscillations can be interpreted as an indication of spin-polaron formation (see, e.g., [Moeller2013]).
Several aspects of the relaxation dynamics in the regime of $J\ll t_{0}$ can be understood by considering the relaxation of a two-level system coupled to a generic bosonic bath, for which there is a substantial amount of literature [seke1983, golosov1999, berman2010, crowley2022]. Essentially, the conduction electron’s polarization is transferred in a single spin-flip process by creating an excitation in the spin bath, thus preventing the electronic kinetic energy from further relaxation. After a short coherent dynamics, there is a Markovian regime with an exponential decay with a time scale $J^{2}/t_{0}$. For $J/t_{0}\lesssim 1.5$, we find an excellent agreement between numerical data and these approximate analytical results.
@@ III.1 Numerical illustration of relaxation dynamics
The initial state with a complete parallel alignment of all localized spins and antiparallel spin orientation of the conduction electron belongs to the subspace $S^{z}_{\text{total}}=\frac{1}{2}(L-1)$. In principle, this coincides with the total spin quantum number of the ground state of the Kondo-lattice model with one itinerant electron. However, in the ground state, the lowest energy is obtained in the momentum sector $K=0$ when the conduction electron has the momentum $k_{0}=0$, implying that our initial state is well-separated from the ground state [sigrist91]. For small $J/t_{0}$ (in the range $0.1<J/t_{0}<2.0$), the initial state becomes essentially orthogonal to the ground state, with its full spectral weight residing in the excited states, as illustrated in App. A.
@@ III.1.1 Regime of $J\ll 2t_{0}$
Based on these observations, we conclude that, in the limit of $J\ll 2t_{0}$, the conduction electron undergoes a scattering process with four components: the electron acquires a backward-propagating ($bp$) and a forward-propagating ($fp$) component, and the same for the emitted magnon [mitrofanov2020, tramsen2021]. After this scattering process, the $z$-component of the electron spin is spin-down and $k_{bp}=-k_{fp}$. Due to momentum conservation, one magnon is excited with weight at the opposite momenta $q^{m}_{bp}=-q^{m}_{fp}$ (up to umklapp scattering).
@@ III.1.2 Two-level picture for $J/t_{0}\ll 1$
For such a situation, standard approximations (see App. B) yield a Markovian behavior and hence an exponential decay is found [prokof2000], in agreement with the exact solution [Henning2012]. We extract the decay constant from a fit to the numerical data, resulting in these functional dependencies for the expectation values:
@@
In Fig. 2, we compare the computational results (solid lines) with the TLS results (dashed lines), where one can observe a good agreement for all operators up to $J/t_{0}\lesssim 1.5$. Note that as $J/t_{0}$ increases, oscillations appear. These correspond to magnon reabsorptions and reemissions not captured in the Markovian approximation. Further, the very short-time window is not captured either since there, non-Markovian processes are always present (see, e.g., [Buesser2014, deVega2017]).
@@ III.1.3 Energy transfer and crossover regime $J\sim 2t_{0}$
For small $J/t_{0}$, a complete magnetization transfer takes place on time scales proportional to $4t_{0}/J^{2}$. At the same time, the kinetic energy of the electron cannot be further reduced after this process since the energy transfer is tied to magnetization transfer. In the small $J/t_{0}$ regime, the energy transferred $\Delta E$ to the spin system is proportional to the exchange coupling. However, as shown in Fig. 5, this linear dependence breaks down when $J/t_{0}\gtrsim 1$. In this intermediate coupling regime, the momentum transfer approaches $\Delta k\approx\pi/2$ as $J/t_{0}\approx 2$. Thus, the initial state couples to the region of the maximum density of states in the dispersion. Therefore, as mentioned before, additional effects beyond the simple TLS picture appear, such as repeated absorption and emission of the magnon, the van-Hove singularity at the band bottom becomes increasingly relevant [Henning2012], and the overlap of the initial state becomes significant with more than two many-body eigenstates. Consequently, the TLS description breaks down and fails to capture the relaxation dynamics for $J\gtrsim t_{0}$.
@@ III.1.4 Regime of $J\gg 2t_{0}$
with fitting parameters $a=0.75$ and $b=0.72$. The Gaussian decay indicates a non-Markovian dynamics [deVega2017, coish2006], which is represented as a black dashed line in Fig. 6. Such a Gaussian decay is also observed in central spin models, where it arises due to a contribution of many individual local environments of localized spins seeing a fluctuating mean-field due the interaction with the central spin [Urbaszek2013, Khaetskii2002, coish2006, Merkulov2002].
In the long-time regime ($t\gg 1/t_{0}$), shown in the inset of Fig. 6 using a log-log scale, the envelope follows a straight line (red dashed line), suggesting a power-law decay proportional to $t^{-3/2}$ due to the van-Hove singularity at the band bottom, in agreement with the exact analytical solution [Henning2012].